X 5 Calculating a bond price - Excel ? X = INSERT - Sign In FILE HOME PAGE LAYOUT FORMULAS DATA REVIEW VIEW Calibri 11 AA - APLEASE WRITE THE ANSWER AS A EXCEL FORMULAAll answers must be entered as a formula Click OK to begin. OK
Answer #1

The bond price can be found using the excel function PV

PV(rate, nper, pmt, [fv], [type])

Cell D17 =PV(D11/2, (D7 - D6)/365*2, D12*D8/2, D12)

rate = 3.8%/2

rate = 0.019

nper = 13 * 2

nper = 26 semi-annual coupon payments

pmt = 2000 * 0.049/2

pmt = 49

FV = 2000

=PV(0.019,26, 49,2000)

-$2,224.043854

(Ignore the negative sign)

Bond price = $2,224.043854

Can you please upvote? Thank You :-)